# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@ Appendix A Appendix
$c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+R^{2})=c_{1}\cdot(\mathfrak{r}^{2}\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}+\mathfrak{r}^{2})\leq c_{1}\cdot(\operatorname{L}+1)\cdot\mathfrak{r}^{2}.$
Thus, for any $\bar{c}\in\mathbb{R}_{\geq 1}$, we can choose $c_{1}=c_{1}(\bar{c},\operatorname{L},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{\geq 1}$ sufficiently large, that
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot(\operatorname{L}+1)\cdot\mathfrak{r}^{2}\right)\leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq 2(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+R^{2})\right\}\right)\leq c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$
@@
$\mathbb{E}_{\theta_{0}}^{N}\left[\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|^{2}_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right]$ $\leq c_{3}\cdot\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{3}\cdot\left((\operatorname{L}+1)\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{4}\cdot\mathfrak{r}_{N}^{2}$
for some constant $c_{4}=c_{4}(\operatorname{L},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{>0}$. This proves the claim.
Proof of 8 (Corollary 2.10).
In Corollary 2.7, we have seen that for each $\bar{c}\in\mathbb{R}_{\geq 1}$ we have
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot\mathfrak{r}^{2}\right)\leq c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}),$
for a constant $c_{1}=c_{1}(\bar{c},,\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{\geq 1}$ large enough, which is equivalent with
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}_{N}^{2}\left\|\hat{\theta}_{\textsc{MAP}}\right\|^{2}_{H^{\alpha}(\mathcal{O},W)}\leq c_{1}\mathfrak{r}_{N}^{2}\right)\geq 1-c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$
Thus, to prove the claim, it suffices to restrict ourselves on the measurable event
$T_{N}:=\left\{\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq c_{1}\mathfrak{r}_{N}^{2},\,\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\sqrt{c_{1}}\right\}.$
As $\theta_{\mathfrak{o}}\in H^{\alpha}(\mathcal{O},W)$ is fixed, there exists some constant $\operatorname{L}=\operatorname{L}(\theta_{\mathfrak{o}})\in\mathbb{R}_{>0}$ with $\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}$ and thus some $c_{2}=c_{2}(c_{1},\operatorname{L})\in\mathbb{R}_{>0}$, such that on the event $T_{N}$, we have
$\left\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq c_{2}\cdot\mathfrak{r}_{N}^{2}\text{ as well as }\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}+\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq c_{2}.$
By an application of C7) in 2.9, we have on the event $T_{N}$
$\left\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c(C_{\operatorname{M}},c_{2})\cdot\mathfrak{r}_{N}^{\tau},$
which shows the claim.
   Proofs of Section 3
@@
   Darcy Problem - Proofs
Let $f\in C^{1}(\bar{\mathcal{O}})$ with $f\geq f_{\min}\in\mathbb{R}_{>0}$. We define the differential operator
@@
where $\nabla$ denotes the usual gradient and $\operatorname{div}$ the divergence operator. Refering on Evans [2010], it is well known from elliptic PDE theory, that $L^{D}$ is a linear and bounded operator, which posses a linear inverse, denoted by
@@
Utilizing the Feynman-Kac formula, Lemma 20 in Nickl et al. [2020] demonstrates that
$\left\|V^{D}_{f}[h]\right\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|h\right\|_{\infty},$ (A.8)
for all $h\in C^{\eta}(\mathcal{O})$, $\eta\in\mathbb{R}_{>0}$, if additionally $f\in C^{2}(\mathcal{O})$. Keeping the notation of Section 3 in mind, we now verify that the forward map $\mathscr{G}^{D}$ satisfies the conditions C1) - C3) formulated in 2.3. As already suggested in the previous section, C3) is the cruicial part for the Darcy problem, as it fails to satisfy the assumption with $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
@@ 1. C1)
    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq C_{\operatorname{Lip},2}\cdot(1+\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{4}\lor\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O})}^{4})\cdot\left\|\theta_{1}-\theta_{2}\right\|_{(H^{1}(\mathcal{O}))^{*}}$
    with some universal constant $C_{\operatorname{Lip},2}\in\mathbb{R}_{>0}$. The proof can be found in Nickl et al. [2020].
  2. C1)
    Uniform boundedness: As described in Section 3, we identify with each $\theta\in H^{\alpha}_{c}(\mathcal{O})$ some $f=f_{\theta}\in\mathcal{F}$, such that $\mathscr{G}^{D}(\theta)=G^{D}(f_{\theta})=V_{f_{\theta}}^{D}[g]$. Thus, with Eq. A.8, we obtain
    $\sup_{\theta\in H^{\alpha}_{c}(\mathcal{O}),x\in\mathcal{O}}\left\{|\mathscr{G}^{D}(\theta)(x)|\right\}$ $=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|G^{D}(f)(x)|\right\}=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|V_{f}^{D}[g](x)|\right\}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|g\right\|_{\infty}.$
  3. C1)
@@
    where the right-hand side defines a function in $C^{\nu}(\mathcal{O})$ for some small $\nu\in\mathbb{R}_{>0}$. In other words, applying the linear inverse $V_{f_{\theta_{1}}}^{D}$ on both sides of the equation, we obtain
    $u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}=V_{f}^{D}[\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})].$
    Thus, by an applycation of Eq. A.8
    $\left\|u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}\right\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}.$
    Moreover, by an applycation of the product rule
    $\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}$ $\leq\left\|\nabla(f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}}\right\|_{\infty}+\left\|(f_{\theta_{2}}-f_{\theta_{1}})\Delta u_{f_{\theta_{2}}}\right\|_{\infty}$
    $\leq\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}\cdot\left(\left\|u_{f_{\theta_{2}}}\right\|_{C^{1}(\mathcal{O})}+\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}\right).$
@@
    $\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}$ $=\left\|f_{\theta_{2}}^{-1}\cdot(g-\nabla u_{f_{\theta_{2}}}\cdot\nabla f_{\theta_{2}})\right\|_{\infty}\leq\left\|f_{\theta_{2}}^{-1}\right\|_{\infty}\left(\left\|g\right\|_{\infty}+\left\|u_{f_{\theta_{2}}}\right\|_{C^{1}(\mathcal{O})}\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}\right).$
@@
[…]
    Thus, elementary computation show that
    $\left\|\Delta u_{f_{\theta_{2}}}\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})(1+\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3}).$
    Hence, by an further application of Lemma 22 in Nickl et al. [2020], we have
    $\left\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3})\cdot\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}.$
    Overall, we have
    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}^{3}\lor\left\|f_{\theta_{2}}\right\|_{C^{1}(\mathcal{O})}^{3})\cdot\left\|f_{\theta_{2}}-f_{\theta_{1}}\right\|_{C^{1}(\mathcal{O})}.$
    By an application of Lemma 29 in Nickl et al. [2020] and Lemma A.10, we obtain finally
    $\left\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\right\|_{\infty}\leq c(f_{\min},\left\|g\right\|_{\infty},d,\mathcal{O})\cdot(1+\left\|\theta_{1}\right\|_{C^{1}(\mathcal{O})}^{4}\lor\left\|\theta_{2}\right\|_{C^{1}(\mathcal{O})}^{4})\cdot\left\|\theta_{2}-\theta_{1}\right\|_{C^{1}(\mathcal{O})}.$
    Thus, the forward map $\mathscr{G}^{D}$ satisfies C3) in 2.3 with $\gamma_{2}=4$, and $\eta_{1}=\eta_{2}=1$.
@@
$\left\|u\right\|_{H^{\beta}(\mathcal{O})}\leq c(\alpha,\beta,\mathcal{O})\cdot\left\|u\right\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot\left\|u\right\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{\beta}{\alpha+1}}.$ (A.9)

### v3

@@ 4.1 Proofs of section 2
$c_{1}\cdot(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+R^{2})=c_{1}\cdot(\mathfrak{r}^{2}\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)}^{2}+\mathfrak{r}^{2})\leq c_{1}\cdot(\operatorname{L}^{2}+1)\cdot\mathfrak{r}^{2}.$
Thus, for any $\bar{c}\in\mathbb{R}_{\geq 1}$, we can choose $c_{1}=c_{1}(\texttt{P},L)\in\mathbb{R}_{\geq 1}$ sufficiently large, that
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq c_{1}\cdot(\operatorname{L}^{2}+1)\cdot\mathfrak{r}^{2}\right)\\ \leq\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\hat{\theta}_{\textsc{MAP}},\theta_{\mathfrak{o}})\geq 2(\mathfrak{d}_{\mathfrak{r}}^{2}(\tilde{\theta},\theta_{\mathfrak{o}})+R^{2})\right)\leq c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$ (25)
@@
$\mathbb{E}_{\theta_{0}}^{N}\left[\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|^{2}_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\right]\leq c_{3}\left(\mathfrak{d}_{\mathfrak{r}}^{2}(\theta_{\mathfrak{o}},\theta_{\mathfrak{o}})+\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{3}\left((\operatorname{L}^{2}+1)\mathfrak{r}_{N}^{2}+\frac{1}{N}\right)\leq c_{4}\mathfrak{r}_{N}^{2}$
for some constant $c_{4}=c_{4}(\operatorname{L},\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W})\in\mathbb{R}_{>0}$. Taking the supremum over all admissible $\theta_{\mathfrak{o}}$ on both sides shows the claim. ∎
Proof of corollary 2.10.
Firstly, for $c_{1}\in\mathbb{R}_{>0}$ define the event
$A_{N}:=\left\{\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}+\mathfrak{r}_{N}^{2}\|\hat{\theta}_{\textsc{MAP}}\|^{2}_{H^{\alpha}(\mathcal{O},W)}\leq c_{1}\mathfrak{r}_{N}^{2}\right\}.$
By an application of corollary 2.7, we can choose for each $\bar{c}\in\mathbb{R}_{\geq 1}$ the constant $c_{1}=c_{1}(\bar{c},,\alpha,\eta_{1},\eta_{2},\kappa,\gamma_{1},\gamma_{2},C_{\operatorname{Lip},2},C_{\operatorname{Lip},\infty},C_{\mathscr{G},\operatorname{B}},d,d_{W},\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)})\in\mathbb{R}_{\geq 1}$ large enough to obatin
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(A_{N}\right)\geq 1-c(d_{V})\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2}).$
We observe that
$A_{N}\subset B_{N}:=\left\{\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq c_{1}\mathfrak{r}_{N}^{2},\,\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}\leq\sqrt{c_{1}}\right\}.$
As $\theta_{\mathfrak{o}}\in H^{\alpha}(\mathcal{O},W)$ is fixed, there exists some constant $\operatorname{L}=\operatorname{L}(\theta_{\mathfrak{o}})\in\mathbb{R}_{>0}$ with $\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{L}$ and thus on the event $B_{N}$, we have
$\|\mathscr{G}(\hat{\theta}_{\textsc{MAP}})-\mathscr{G}(\theta_{\mathfrak{o}})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}^{2}\leq c_{1}\cdot\mathfrak{r}_{N}^{2}\text{ as well as }\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}+\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)}\leq\sqrt{c_{1}}+L.$
By applying the inverse continuity modulus formulated in 2.9, we have on the event $B_{N}$
$\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c(C_{\operatorname{M}},c_{1})\cdot\mathfrak{r}_{N}^{\tau}.$
With the estimations above, the claim follows.
4.2 Proofs of section 3
@@
where $\nabla$ denotes the usual gradient and $\operatorname{div}$ the divergence operator. From classical elliptic PDE theory (see for instance Evans (2010) Theorem 4 and the Remark after it in Chapter 6.3), it is well-known, that $L^{D}$ is a linear and bounded operator, which possess a linear inverse, we denote by
@@
Utilizing the corresponding Feynman-Kac formula, Lemma 20 in Nickl et al. (2020) demonstrates that
$\|V^{D}_{f}[h]\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\|h\|_{\infty}$ (26)
holds for all $h\in C^{\eta}(\mathcal{O})$, $\eta\in\mathbb{R}_{>0}$, if additionally $f\in C^{2}(\mathcal{O})$. Keeping the notation of section 3 in mind, we now verify that the forward map $\mathscr{G}^{D}$ satisfies the conditions C1) - C3) formulated in 2.3. The cruicial part is to verify condition C3) for some $\eta_{1}$, $\eta_{2}$ and $\gamma_{2}$.
@@ 1. C1)
    $\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\|_{\mathbb{L}^{2}(\mathcal{O})}\leq C_{\operatorname{Lip},2}\cdot(1+\|\theta_{1}\|_{H^{\alpha}(\mathcal{O})}^{4}\lor\|\theta_{2}\|_{H^{\alpha}(\mathcal{O})}^{4})\cdot\|\theta_{1}-\theta_{2}\|_{(H^{1}(\mathcal{O}))^{*}}$
    with some universal constant $C_{\operatorname{Lip},2}\in\mathbb{R}_{>0}$. The proof can be found in Nickl et al. (2020).
  2. C2)
    Uniform boundedness: As described in section 3, we identify with each $\theta\in H^{\alpha}_{c}(\mathcal{O})$ some $f=f_{\theta}=\Psi(\theta)\in\mathcal{F}$, such that $\mathscr{G}^{D}(\theta)=G^{D}(f_{\theta})=V_{f_{\theta}}^{D}[g]$. Thus, with eq. 26, we obtain
    $\sup_{\theta\in H^{\alpha}_{c}(\mathcal{O}),x\in\mathcal{O}}\left\{|\mathscr{G}^{D}(\theta)(x)|\right\}$ $=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|G^{D}(f)(x)|\right\}$
    $=\sup_{f\in\mathcal{F},x\in\mathcal{O}}\left\{|V_{f}^{D}[g](x)|\right\}\leq c(d,\mathcal{O},f_{\min})\cdot\|g\|_{\infty}.$
  3. C3)
@@
    where the right-hand side defines a function in $C^{\nu}(\mathcal{O})$ for some small $\nu\in\mathbb{R}_{>0}$. Thus, applying the linear inverse $V_{f_{\theta_{1}}}^{D}$ on both sides of the equation, we obtain
    $u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}=V_{f}^{D}[\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})],$
    and hence, by an application of eq. 26
    $\|u_{f_{\theta_{1}}}-u_{f_{\theta_{2}}}\|_{\infty}\leq c(d,\mathcal{O},f_{\min})\cdot\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\|_{\infty}.$
    Utilizing the product rule, the right-hand side can be further upper bounded by
    $\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\|_{\infty}$ $\leq\|\nabla(f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}}\|_{\infty}+\|(f_{\theta_{2}}-f_{\theta_{1}})\Delta u_{f_{\theta_{2}}}\|_{\infty}$
    $\leq\|f_{\theta_{2}}-f_{\theta_{1}}\|_{C^{1}(\mathcal{O})}\cdot\left(\|u_{f_{\theta_{2}}}\|_{C^{1}(\mathcal{O})}+\|\Delta u_{f_{\theta_{2}}}\|_{\infty}\right).$
@@
    $\|\Delta u_{f_{\theta_{2}}}\|_{\infty}$ $=\|f_{\theta_{2}}^{-1}\cdot(g-\nabla u_{f_{\theta_{2}}}\cdot\nabla f_{\theta_{2}})\|_{\infty}$
    $\leq\|f_{\theta_{2}}^{-1}\|_{\infty}\cdot\left(\|g\|_{\infty}+\|u_{f_{\theta_{2}}}\|_{C^{1}(\mathcal{O})}\|f_{\theta_{2}}\|_{C^{1}(\mathcal{O})}\right)$
@@
    and hence
    $\|\Delta u_{f_{\theta_{2}}}\|_{\infty}\leq c(f_{\min},\|g\|_{\infty},d,\mathcal{O})\cdot(1+\|f_{\theta_{2}}\|_{C^{1}(\mathcal{O})}^{3}).$
    Another application of Lemma 22 in Nickl et al. (2020) yields
    $\|\operatorname{div}((f_{\theta_{2}}-f_{\theta_{1}})\cdot\nabla u_{f_{\theta_{2}}})\|_{\infty}\leq c(f_{\min},\|g\|_{\infty},d,\mathcal{O})\cdot(1+\|f_{\theta_{2}}\|_{C^{1}(\mathcal{O})}^{3})\cdot\|f_{\theta_{2}}-f_{\theta_{1}}\|_{C^{1}(\mathcal{O})},$
    which leads overall to
    $\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\|_{\infty}\\ \leq c(f_{\min},\|g\|_{\infty},d,\mathcal{O})\cdot(1+\|f_{\theta_{1}}\|_{C^{1}(\mathcal{O})}^{3}\lor\|f_{\theta_{2}}\|_{C^{1}(\mathcal{O})}^{3})\|f_{\theta_{2}}-f_{\theta_{1}}\|_{C^{1}(\mathcal{O})}.$
    Now utilizing the properties of the regular link function $\Psi$ formulated in lemma 4.10 and Lemma 29 in Nickl et al. (2020), it follows
    $\|\mathscr{G}^{D}(\theta_{1})-\mathscr{G}^{D}(\theta_{2})\|_{\infty}\\ \leq c(f_{\min},\|g\|_{\infty},d,\mathcal{O})\cdot(1+\|\theta_{1}\|_{C^{1}(\mathcal{O})}^{4}\lor\|\theta_{2}\|_{C^{1}(\mathcal{O})}^{4})\cdot\|\theta_{2}-\theta_{1}\|_{C^{1}(\mathcal{O})}.$
    Hence, the forward map $\mathscr{G}^{D}$ satisfies C3) in 2.3 with $\gamma_{2}=4$, and $\eta_{1}=\eta_{2}=1$.
@@
$\|u\|_{H^{\beta}(\mathcal{O})}\leq c(\alpha,\beta,\mathcal{O})\cdot\|u\|_{\mathbb{L}^{2}(\mathcal{O})}^{\frac{\alpha+1-\beta}{\alpha+1}}\cdot\|u\|_{H^{\alpha+1}(\mathcal{O})}^{\frac{\beta}{\alpha+1}},$ (27)
which is proven for instance in Theorem 1 in Section 4.3.1 in Triebel (1978) observing that $H^{\beta}(\mathcal{O})=B_{22}^{\beta}(\mathcal{O})$.